OpenStudy (anonymous):

Find a polynomial function with the given zeros, 0,0,0,sqrt6, -sqrt6

Parth (parthkohli):

\[(x-0)(x-0)(x-0)(x-\sqrt{6})(x+\sqrt6)\]

Parth (parthkohli):

Of course, \((x-0)(x-0)(x-0)=x^3\)

OpenStudy (anonymous):

So would the answer be x^5-6x^3?

OpenStudy (whpalmer4):

Yes, that is correct.

OpenStudy (anonymous):

Yeah,,, thanks

OpenStudy (whpalmer4):

Here's a plot showing the axis crossings at the roots of odd multiplicity

OpenStudy (whpalmer4):

even multiplicity gets you a touching of the axis, but not a crossing

OpenStudy (whpalmer4):

vertical lines indicate \(x=\pm\sqrt{6}\)